BOUND AND SCATTERING STATE SOLUTIONS OF SCHRODINGER EQUATION FOR ASYMMETRIC WOODS SAXON POTENTIAL

N. CANDEMIR, Physics Department, Science Faculty, Anadolu University, Eskişehir, 26470, Turkey.

The one-dimensional time-independent Schrödinger equation is solved for asymmetric Woods-Saxon potential. The reflection and transmission coefficients and bound state solutions are obtained in terms of hypergeometric functions. Some useful figures are plotted to show the accuracy of the obtained results.